OpenStudy (anonymous):

Use a transformation to solve the equation. p + (–26) = –13.8 p = 39.8 p = –12.2 p = 12.2 p = –39.8

OpenStudy (welshfella):

p + (-26) = -13.8 p - 26 = -13.8 now add 26 to both sides and you have the solution

OpenStudy (anonymous):

so it would be p= 39.8

OpenStudy (welshfella):

no -13.8 + 26 = 26 - 13.8 = ?

OpenStudy (welshfella):

its negative 13.8 not 13.8

OpenStudy (welshfella):

answer is 26 - 13.8